OpenStudy (anonymous):

solve the equation for m: a=2(m-w)

OpenStudy (anonymous):

divide the 2 over to the other side\[a/2=m-w\]add w over to the other side\[m= a/2+w\]

OpenStudy (anonymous):

helpful?? @heartbreak :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

the answer would be m=2(a+w)

OpenStudy (anonymous):

no your answer is m=a/2 + w

OpenStudy (anonymous):

@heartbreak